# Closed-Form Interpretation of Neural Network Classifiers with Symbolic Regression Gradients

### v1

@@ Abstract
I introduce a unified framework for interpreting neural network classifiers tailored toward automated scientific discovery. In contrast to neural network-based regression, for classification, it is in general impossible to find a one-to-one mapping from the neural network to a symbolic equation even if the neural network itself bases its classification on a quantity that can be written as a closed-form equation. In this paper, I embed a trained neural network into an equivalence class of classifying functions that base their decisions on the same quantity. I interpret neural networks by finding an intersection between this equivalence class and human-readable equations defined by the search space of symbolic regression. The approach is not limited to classifiers or full neural networks and can be applied to arbitrary neurons in hidden layers or latent spaces or to simplify the process of interpreting neural network regressors.
@@
1 TL;DR
What do you do in practice? To interpret an artificial neural network, I fit the gradients of a symbolic regression model to the gradients of the output neuron of the neural network.
@@ 2 Introduction
Artificial neural networks have become the most powerful tool in machine learning applications in a wide range of applications such as image classification, speech recognition, or natural language processing. In many fields, artificial neural networks have achieved better than human performance, and the number of these fields is consistently growing. The larger and more powerful the artificial neural networks become the more elusive is their decision making to us humans. The black-box nature of these models makes it very challenging to comprehend their decision-making processes.
Interpreting artificial neural networks is a central endeavor in many applications. As artificial neural networks are entering domains where a detailed understanding of the decision-making process is crucial, it becomes increasingly important to interpret their inner workings. These fields contain safety-critical, medical [1, 2] applications where human lives could be in danger. The legislative domains [3, 4] in which each decision needs to be supported by law. Self-driving cars are in danger of causing accidents because of a lack of understanding of the underlying machine learning algorithms [5]. Further, in many scientific applications [6] the primary utility of a model is to help understand an underlying phenomenon much more than making accurate predictions.
@@
To address the problem of interpreting neural network classifiers, it is important to acknowledge that concepts encoded in an artificial neural network are stored in a highly convoluted and elusive manner. A neural network does not encode them in a human-readable form, but through complex interactions of thousands if not millions of neurons/perceptrons. In other words, even if the neural network learns a concept that could be put into a human-readable form, a neural network hides this concept through complex and highly nonlinear transformations.
This observation leads to the initial idea behind the interpretation framework proposed in this manuscript. Is it possible to extract a human-readable concept $g$ that is learned by a neural network $F$ by getting rid of the elusive transformation $\phi$ that conceals it.
@@
The current paper provides an answer to this question by embedding the neural network $F$ into an equivalence class of functions that learn the same concept. A closed-form expression of this concept can be recovered by finding the intersection of this equivalence class with the space of human-readable functions defined through a symbolic regression search space.
2.1 Outline
This article is structured in the following manner: After exploring the most closely related prior research on neural network interpretability in Section 3:Prior Work, I describe the formal and mathematical framework that is necessary for the interpretation algorithm in Section 4: Theoretical Framework. Next, I explain the computational implementation of the interpretation algorithm in Section 5:Interpretation Algorithm. Finally, I perform experiments demonstrating interpretation results in Section 6:Experiments. The findings and potential follow-up research are summarized in Section 7: Conclusion.
3 Prior Work
To the best of my knowledge, there are no successful approaches that manage to find closed-form expressions of the high-level concepts learned by neural network classifiers beyond reducing the problem to perceptrons.
@@ 3 Prior Work
On the other hand, it is possible to directly employ symbolic classification approaches to obtain a symbolic solution to a classification problem. This can be done by performing symbolic regression with margin or hinge losses.
Common symbolic regression libraries that mostly employ genetic algorithms include Eureqa, which was employed in [12], Operon C++[26], PySINDy [27], Feyn[28], Gene-pool Optimal Mixing Evolutionary Algorithm [29], GPLearn [30] and PySR[31]. There are also symbolic regression algorithms that employ neural networks, for example, EQL with embedded symbolic layers[32, 33], deep symbolic regression uses recurrent neural networks [34], and symbolic regression with transformers[35]. Further, AI Feynman uses neural networks to simplify expressions [36]. An overview of interpretable scientific discovery with symbolic Regression can be found in[37]
4 Theoretical Framework
Figure 1: An artificial neural network is a connected graph consisting of nodes representing neurons and weighted connections between them. A neural network predicts an approximate target $F(\mathbf{x})=\hat{y}\approx y$ here in the context of binary classification. Removing the final sigmoid activation function allows the extraction of the latent model $f$ from the full neural network $F$ for easier interpretability.
4.1 Artificial Neural Network
An artificial neural network fig. 1 is a connected graph containing nodes representing neurons otherwise known as perceptrons and connections representing weighted inputs that are supplied to these nodes. A basic feedforward neural network can be represented mathematically as follows:
A neural network $F$ is applied to an input data point $\mathbf{a}^{(0)}=\mathbf{x}\in\mathbb{R}^{n}$ represented through an embedding in an $n$-dimensional space of real numbers. The input gets processed through a number $l$ of layers $L$ containing weights matrices $\mathbf{W}\in\mathbb{R}^{m_{l}\times m_{l-1}}$, biases vectors $\mathbf{b}\in\mathbb{R}^{m_{l}}$ and non-linear activation functions $\sigma$
$\mathbf{a}^{(l)}=L^{l}(\mathbf{a}^{(l-1)})=\sigma^{(l)}(\mathbf{W}^{(l)}\mathbf{a}^{(l-1)}+\mathbf{b}^{(l)})$ (2)
to produce some output
The parameters of an artificial neural network are trained on a data set $(X,Y)$ representing the data manifold $X\subset D\subset\mathbb{R}^{n}$ through gradient descent/backpropagation to minimize an objective/loss/error function $L(F(X),Y)$ such that $F(X)\approx Y$ on all data points on the data manifold.
@@ 4.2 Interpreting an Artificial Neural Network
Interpreting an artificial neural network fig. 1 means formulating a mapping between an abstract mechanism or encoding into a domain that a human can understand. Human understanding is a vaguely defined concept since human capabilities range from simply perceiving common objects in the real world to expert knowledge in their respective field. I suggest categorizing interpretations along the following four different attributes, which is more fine-grained than the categories proposed in [10]:
@@
C) Verifying vs. Discover: A machine learning practitioner might have a set of hypotheses of concepts in mind that a neural network might learn in order to solve a given task. It is often easy to verify or falsify if these concepts are included in the features learned by an artificial neural network. However, discovering an unknown concept without a list of potential hypotheses is in general very hard.
@@
Figure 2: The interpretation framework is based on finding the intersection between human-readable functions and the equivalence class (6) of functions that contain the same information as an output neuron of the neural network. The space of human-readable functions can be defined through a symbolic regression search space with elementary functions and complexity that matches the user’s knowledge.
@@ 4.3 Equivalence Class of Functions Containing the Same Information
4.3.1 Definition
The goal of the interpretation framework outlined in this manuscript is to find the intersection of the equivalence class of functions that contain the same information as the output neuron (or any other neuron) and the set of human-readable equations, see fig. 2.
@@ 4.3.1 Definition
If $f(\mathbf{x})$ is the output of an artificial neural network, this defines an equivalence relation between all functions that base their decision on the same quantity $g(\mathbf{x})$. Several neural networks might encode the same function $f$ because it does not matter how the output is calculated mechanistically. All realizations of a neural network that is symmetric under permutations or other changes in its weights would still be functionally equivalent and thus the same function within the equivalence class $\tilde{H}_{g}$.
@@
Using this property, I define the equivalence set $H_{g}\subseteq\tilde{H}_{g}$, which will later form the basis for the neural network interpretation algorithm:
@@
It remains to show that by using the above definitions: $H_{g}=\tilde{H}_{g}$. Hence, one needs to show that for each function $f\in H_{g}\ \exists\phi:f(\mathbf{x})=\phi(g(\mathbf{x}))$. It is possible to explicitly construct the function $\phi$ that maps between $f$ and $g$. Defining $\phi^{\prime}$ through (5) leads to an integrable $\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})=\nabla f(\mathbf{x})$ because a) the images of $f(D)$ and $g(D)$ are compact, thus $\phi^{\prime}$ maps between compact subsets of $\mathbb{R}$ and b) $\phi^{\prime}$ is continuous. For any simply connected $D\subset\mathbb{R}^{n}$ I can define the $C^{1}$-curve $\mathbf{x}:[t_{0},t_{1}]\rightarrow D$.
@@ 4.3.2 Assumptions
In a practical machine learning application not all assumptions from the prior section that ensure $H_{g}=\tilde{H}_{g}$ hold true. However, even then $H_{g}\approx\tilde{H}_{g}$ can provide a good approximation that allows for a retrieval of the function that a neuron encodes.
@@
4.4 Symbolic Regression Search Space
4.4.1 Human-Readable Functions
@@
Throughout this paper, I base the set of human-readable functions on symbolic regression trees, see fig. 3. These trees can be constructed to contain a user-defined set of elementary operations. Symbolic trees can also be associated with a complexity measure that increases the larger the tree grows leading to several solutions along the Pareto front. Hence, it is possible to address human readability according to aspects 1) and 2). Within the scope of this paper, it is impossible to address the question of the context that gives the equation meaning, as discussed in 3). In recent years, progress in artificial scientific discovery has made it possible to train neural networks to optimize loss functions based on concepts instead of regression targets [18, 20, 22, 21]. The interpretation framework developed in the current manuscript is uniquely suitable to address the interpretation in these cases.
Figure 3: Symbolic regression tree that represents a mathematical function. A tree consists of connected nodes containing variables, numeric parameters, unary and binary operators. Symbolic regression is performed by a genetic algorithm that modifies, evolves and adds nodes to optimize an objective function on some underlying training data.
4.4.2 Symbolic Regression Algorithm
In order to computationally search through the space of human-readable functions and determine the intersection with the equivalence set described in the previous section, it is important to build upon a suitable symbolic regression algorithm.
In principle, a simple way of searching through the space of possible functions could be optimizing a linear combination of user-defined elementary functions. A much better approach can be taken by utilizing evolutionary or genetic algorithms. These algorithms build trees of connected nodes that represent a mathematical function, see fig. 3. Each node represents input variables, numeric parameters, as well as unary and binary operators. A genetic algorithm modifies, evolves, and adds nodes to optimize an objective function on some underlying training data.
@@ 4.4.2 Symbolic Regression Algorithm
5 Interpretation Algorithm
In this section, I discuss how to interpret a neural network classifier $F$ in practice by finding a representative symbolic function that lives on the intersection between human-readable functions and all functions that contain the same information as a trained neural network, see fig. 2. While the description here focuses on binary classifiers, it is possible to apply the framework to interpret multi-class classifiers or even neurons contained in the hidden layers of neural networks. The interpretation framework contains three algorithmic steps: 1) training an artificial neural network 1, 2) extracting gradients of the output neuron with respect to the input 2 and 3) performing symbolic regression to find a symbolic function whose gradients approximate the gradients from the desired neuron of an artificial neural network 3.
Figure 4: (a) Two class data of Experiment 1 separated by decision boundary $g(\mathbf{x})=x_{1}^{2}+2x_{2}^{2}=1$. A neural network $F$ is trained to classify the data. Afterward, a symbolic model $T$ is trained to reproduce the normalized gradients of $F$ which coincide with the normalized gradients of function that defines the decision boundary $g$. (b) Empirical correlation between true function $g$ and the neural network $F$. Removing the sigmoid activation function from $F$ defines the latent model $f$ which has an almost linear correlation with $g$. However, this correlation is not linear and defines the function $f=\phi(g)$ with which I ascertain the equivalence relation $f\sim g$ assuring that $F$ and $f$ contain the same information as $g$ and thus $F,f\in\tilde{H}_{g}$ according to eq.(4).
5.1 Training a Neural Network For Binary Classification
The first part of the interpretation procedure involves training an artificial neural network for binary classification on a two-class data set, see algorithm 1, to approximate the decision boundary to separate the two classes, see fig. 4(a). After successful training I obtain the full model $F$ and the latent model $f$ by removing the final sigmoid activation function, which has a cleaner correlation to a closed-form expression of the decision boundary $g$, see fig. 4(b):

### v2

# Closed-Form Interpretation of Neural Network Classifiers with Symbolic Gradients

@@ Abstract
I introduce a unified framework for finding a closed-form interpretation of any single neuron in an artificial neural network. Using this framework I demonstrate how to interpret neural network classifiers to reveal closed-form expressions of the concepts encoded in their decision boundaries. In contrast to neural network-based regression, for classification, it is in general impossible to express the neural network in the form of a symbolic equation even if the neural network itself bases its classification on a quantity that can be written as a closed-form equation. The interpretation framework is based on embedding trained neural networks into an equivalence class of functions that encode the same concept. I interpret these neural networks by finding an intersection between the equivalence class and human-readable equations defined by a symbolic search space. The approach is not limited to classifiers or full neural networks and can be applied to arbitrary neurons in hidden layers or latent spaces.
@@ 1 Introduction
Artificial neural networks have become the most powerful tool in a wide range of machine learning topics such as image classification, speech recognition, or natural language processing. In many fields, artificial neural networks have achieved better than human performance, and the number of these fields is consistently growing. The larger and more powerful the artificial neural networks become the more elusive is their decision-making to us humans. The black-box nature of these models makes it very challenging to comprehend their decision-making processes.
Interpreting artificial neural networks is a central endeavor in many applications. As artificial neural networks are entering domains where a detailed understanding of the decision-making process is crucial, it becomes increasingly important to interpret their inner workings. These fields contain safety-critical, medical [1, 2] applications where human lives could be in danger or the legislative domains [3, 4] in which each decision needs to be supported by law. Self-driving cars are in danger of causing accidents because of a lack of understanding of the underlying machine learning algorithms [5]. Further, in many scientific applications [6, 7, 8, 9, 10, 11, 12, 13, 14] the primary utility of a model is to help scientists understand an underlying phenomenon much more than making accurate predictions.
@@
To address the problem of interpreting neural network classifiers, it is crucial to acknowledge that concepts encoded in an artificial neural network are stored in a highly convoluted and elusive manner. A neural network does not encode them in a human-readable form, but through complex interactions of thousands if not millions of neurons/perceptrons. In other words, even if the neural network learns a concept that could be put into a human-readable form, a neural network hides this concept through complex and highly nonlinear transformations.
This observation leads to the initial idea behind the interpretation framework proposed in this manuscript. Is it possible to extract a human-readable concept $g$ that is learned by a neural network $F$ by getting rid of the elusive transformation $\phi$ that conceals it?
@@
The current paper provides an answer to this question by embedding the neural network $F$ into an equivalence class of functions that learn the same concept. A closed-form expression of this concept can be recovered by finding the intersection of this equivalence class with the space of human-readable functions defined through a symbolic search space.
To the best of my knowledge, there are no successful approaches that manage to find closed-form expressions of the high-level concepts learned by arbitrary neurons within a neural network. Much research has been done to find symbolic solutions to regression problems of which some can be extended to classification problems.
@@
It was proposed to train a neural network to map the weights of a neural network to a feature vector representing symbolic functions [26].
While their training efficiency and accuracy cannot compete with artificial neural networks, it is possible to directly employ symbolic regression approaches to obtain a symbolic solution that replicates the behaviour of a neural network. Many of these algorithms can be extended to classification problems through margin or hinge loss functions. Common symbolic regression libraries that mostly employ genetic algorithms include Eureqa [20], Operon C++[27], PySINDy [28], Feyn[29], Gene-pool Optimal Mixing Evolutionary Algorithm [30], GPLearn [31] and PySR[32].
There are also symbolic regression algorithms that employ neural networks, for example, EQL with embedded symbolic layers[33, 34, 35], deep symbolic regression uses recurrent neural networks [36], and symbolic regression with transformers[37, 38]. Further, AI Feynman uses neural networks to simplify expressions [39].
An overview of interpretable scientific discovery with symbolic Regression can be found in[40].
It is important to note that the process of replicating the output of a neural network faces a fundamental flaw: there is no guarantee that this procedure captures the same features that a neural network learns to solve a specific problem. Hence, replicating the output of a neural network is different from interpreting a neural network. This observation is demonstrated in section 4.3.
2 Framework Overview
Figure 1: a: An artificial neural network is a connected graph consisting of nodes representing neurons and weighted connections between them. A neural network predicts an approximate target $F(\mathbf{x})=\hat{y}\approx y$ here in the context of binary classification. Removing the final sigmoid activation function allows the extraction of the latent model $f$ from the full neural network $F$ for easier interpretability. b: The interpretation framework is based on finding the intersection between human-readable functions and the equivalence class 4 of functions that contain the same information as an output neuron of the neural network. The space of human-readable functions can be defined through a symbolic search space with elementary functions and complexity that matches the user’s knowledge. This space can be computationally explored by genetic algorithms whose structure is mathematically represented by c: a symbolic tree. A tree consists of connected nodes containing variables, numeric parameters, unary and binary operators. Symbolic search is performed by a genetic algorithm that modifies, evolves and adds nodes to optimize an objective function on some underlying training data.
@@ 2.1 Interpreting an Artificial Neural Network
Interpreting an artificial neural network means formulating a mapping between an abstract mechanism or encoding into a domain that a human can understand. Human understanding is a vaguely defined concept since human capabilities range from simply perceiving common objects in the real world to expert knowledge in their respective field. I suggest categorizing interpretations along the following four different attributes, which is more fine-grained than the categories proposed in [18]:
@@
C) Verify vs. Discover: A machine learning practitioner might have a set of hypotheses of concepts in mind that a neural network might learn in order to solve a given task. It is often easy to verify or falsify if these concepts are included in the features learned by an artificial neural network. However, discovering an unknown concept without a list of potential hypotheses is in general very hard.
@@ 2.2 Equivalence Class of Functions Containing the Same Information
The goal of the interpretation framework outlined in this manuscript is to find the intersection of the equivalence class of functions that contain the same information as the output neuron (or any other neuron) and the set of human-readable equations, see fig. 1b.
@@
If $f(\mathbf{x})$ is the output of an artificial neural network, this defines an equivalence relation between all functions that base their decision on the same quantity $g(\mathbf{x})$. Several different neural networks might encode the same function $f$ because, in the context of this section, it does not matter how the output is calculated mechanistically. All realizations of a neural network that is symmetric under permutations or other changes in its weights would still be functionally equivalent and thus the same function within the equivalence class $\tilde{H}_{g}$.
@@
Using this property, I define the equivalence set $H_{g}$, with $\tilde{H}_{g}\subseteq H_{g}$, which will later form the basis for the neural network interpretation algorithm:
@@
2.3 Equivalence of Equivalence Classes
2.3.1 Proof
Let $g(\mathbf{x}),f(\mathbf{x})\in C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})$ be continuously differentiable functions ($C^{1}(D\subset\mathbb{R}^{n},\mathbb{R})$ is the vector space of differentiable functions from $D$ to $\mathbb{R}$) and $D\subset\mathbb{R}^{n}$ be the data manifold which is required to be compact and simply connected. Then $\tilde{H}_{g}=H_{g}$ where
and
Proof: One can see that for each function $f\in\tilde{H}_{g}\ \phi:\nabla f(\mathbf{x})=\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})$ , hence $\tilde{H}_{g}\subset H_{g}$.
It remains to be shown that for each function $f\in H_{g}\ \exists\phi:f(\mathbf{x})=\phi(g(\mathbf{x}))$. It is possible to explicitly construct the function $\phi$ that maps between $f$ and $g$. Defining $\phi^{\prime}$ through
leads to an integrable $\phi^{\prime}(g(\mathbf{x}))\nabla g(\mathbf{x})=\nabla f(\mathbf{x})$ because a) the images of $f(D)$ and $g(D)$ are compact, thus $\phi^{\prime}$ maps between compact subsets of $\mathbb{R}$ and b) $\phi^{\prime}$ is continuous. For any simply connected $D\subset\mathbb{R}^{n}$ I can define the $C^{1}$-curve $\mathbf{x}:[t_{0},t_{1}]\rightarrow D$, thus a variable transformation within the calculation of the contour integral yields:
@@ 2.3.2 Assumptions
In a practical machine learning applications not all assumptions from the prior section that ensure $H_{g}=\tilde{H}_{g}$ hold true. However, even then $H_{g}\approx\tilde{H}_{g}$ can provide a good approximation that allows for a retrieval of the function that a neuron encodes.
@@
2.4 Extracting Symbolic Concepts Encoded in Neural Networks
A neural network classifier $F$ can be interpreted in practice by finding a representative symbolic function that lives on the intersection between human-readable functions and all functions that contain the same information as a trained neural network, see fig. 1b. For this purpose, I employ a genetic algorithm to find a function represented by a symbolic tree $T$ whose normalized gradients approximate the normalized gradients of a latent neural network model $f$. Here $f$ is defined by removing the final activation function $F(x)=\text{sigmoid}(f(x))$. Through this procedure it is possible to find an element belonging to the equivalence class 4 in human-readable form.
While the experiments here focus on interpreting binary classifiers, it is possible to apply the framework to interpret multi-class classifiers or even neurons contained in the hidden layers of neural networks. The interpretation framework contains three algorithmic steps summarized in algorithm 1: 1) training an artificial neural network, 2) extracting gradients of the output neuron with respect to the input and 3) performing symbolic search to find a symbolic function whose normalized gradients approximate the normalized gradients from the desired neuron of an artificial neural network.
3 Methods
3.1 Artificial Neural Network
An artificial neural network fig. 1a is a connected graph containing nodes representing neurons otherwise known as perceptrons and connections representing weighted inputs that are supplied to these nodes. A basic feedforward neural network can be represented mathematically as follows:
A neural network $F$ is applied to an input data point $\mathbf{a}^{(0)}=\mathbf{x}\in\mathbb{R}^{n}$ represented through an embedding in an $n$-dimensional space of real numbers. The input gets processed through a number $l$ of layers $L$ containing weight matrices $\mathbf{W}\in\mathbb{R}^{m_{l}\times m_{l-1}}$, biases vectors $\mathbf{b}\in\mathbb{R}^{m_{l}}$ and non-linear activation functions $\sigma$
$\mathbf{a}^{(l)}=L^{(l)}(\mathbf{a}^{(l-1)})=\sigma^{(l)}(\mathbf{W}^{(l)}\mathbf{a}^{(l-1)}+\mathbf{b}^{(l)})$ (13)
to produce some output
The parameters of an artificial neural network are trained on a data set $(X,Y)$ representing the data manifold $X\subset D\subset\mathbb{R}^{n}$ through gradient descent/backpropagation to minimize an objective/loss/error function $\mathcal{L}(F(X),Y)$ such that $F(X)\approx Y$ on all data points on the data manifold.
3.2 Symbolic Search
3.2.1 Symbolic Search Space
@@
Throughout this paper, I base the set of human-readable functions on symbolic trees, see fig. 1c. These trees can be constructed to contain a user-defined set of elementary operations. Symbolic trees can also be associated with a complexity measure that increases the larger the tree grows leading to several solutions along the Pareto front. Hence, it is possible to address human readability according to aspects 1) and 2). Within the scope of this paper, it is impossible to address the question of the context that gives the equation meaning, as discussed in 3). In recent years, progress in artificial scientific discovery has made it possible to train neural networks to optimize loss functions based on concepts that imply a context, instead of optimizing for regression targets [9, 11, 13, 12]. The interpretation framework developed in the current manuscript is uniquely suitable to address the interpretation in these cases.
3.2.2 Symbolic Search Algorithm
In order to computationally search through the space of human-readable functions and determine the intersection with the equivalence set described in the previous section, it is convenient to build upon a suitable symbolic regression algorithm based on a backend that performs an efficient search in the space of functions represented by symbolic trees.
In principle, a simple way of searching through the space of possible functions could be optimizing a linear combination of user-defined elementary functions. A much better approach can be taken by utilizing evolutionary or genetic algorithms. These algorithms build trees of connected nodes that represent a mathematical function, see fig. 1c. Each node represents input variables, numeric parameters, as well as unary and binary operators. A genetic algorithm modifies, evolves, and adds nodes to optimize an objective function on some underlying training data.
@@ 3.2.2 Symbolic Search Algorithm
3.3 Interpretation Algorithm
Figure 2: a: Two class data of Experiment 1 separated by decision boundary $g(\mathbf{x})=x_{1}^{2}+2x_{2}^{2}=1$. A neural network $F$ is trained to classify the data. Afterward, a symbolic model $T$ is trained to reproduce the normalized gradients of $F$ which coincide with the normalized gradients of function that defines the decision boundary $g$. b: Empirical correlation between true function $g$ and the neural network $F$. Removing the sigmoid activation function from $F$ defines the latent model $f$ which has an almost linear correlation with $g$. However, this correlation is not linear and defines the function $f=\phi(g)$ with which I ascertain the equivalence relation $f\sim g$ assuring that $F$ and $f$ contain the same information as $g$ and thus $F,f\in\tilde{H}_{g}$ according to eq.2.
3.3.1 Training a Neural Network For Binary Classification
The first part of the interpretation procedure involves training an artificial neural network for binary classification on a two-class data set, see algorithm 1.1, to approximate the decision boundary that separates the two classes, see fig. 2a. After successful training I obtain the full model $F$ and the latent model $f$ by removing the final sigmoid activation function, which has a cleaner correlation to a closed-form expression of the decision boundary $g$, see fig. 2b: